A sperical capacitor has an inner sphere of raduis 10 cm and an outer sphere of raduis 14 cm


1
Expert's answer
2021-05-30T13:23:28-0400
"C=\frac{4πϵϵ _0rR}{R-r}\\C=\frac{4π(32)(8.85\cdot10^{-12})(0.14)(0.1)}{0.14-0.1}=1.25\cdot10^{-9}\ F"